If 55.0 mL of ethanol ) initially at 8.0C is mixed with 55.0 ml of water (density = 1.0g/mL) initially at 28.6C in an insulated beaker, and assuming that no heat is lost, what is the final temperature of the mixture?
Answer #1
Mass of ethanol, m,E = density * volume
                                 = 0.789 g/mL * 55.0 mL
                                 = 43.4 g
Mass of water, m,H2O = 55.0 mL * 1.0 g/mL
                                     = 55.0 g
Given data:
Initial temperature of ethanol is T,E = 8.0 C
Initial temperature of water is T,H2O = 28.6 C
Let, final temperature of mixture be T.
Heat capacity of water is, C,H2O = 4.18 J/gC
Let, heat capacity of ethanol is C,E
As no heat is lost to surroundings, heat gained by ethanol is equal to heat lost by water.
                                   q,E = - q,H2O
                m * C * (T-T,E) = -m * C * (T - T,H2O)
             43.4 g * C * (T - 8.0) = -55.0 g* 4.18 J/gC * (T - 28.6) C
                  43.4 C (T-8.0) = 230 (T-28.6)
Heat capacity of ethanol is not mentioned in the problem, if we know that we can calculate the final temperature.